OpenStudy (anonymous):

What are they asking? The parabola is in position I. y^2= 4x x^2= 6y x^2+ 10y = 0 x/2= y^2 3y = x^2

OpenStudy (anonymous):

@amistre64 Can you help me please?

OpenStudy (anonymous):

these are the options ?

OpenStudy (anonymous):

Yes, but I can choose more than 1

OpenStudy (anonymous):

the answer is first and last i guess

OpenStudy (anonymous):

ok ill try that

OpenStudy (anonymous):

but what exactly are they asking for

OpenStudy (anonymous):

They gave me partial credit for those 2

OpenStudy (anonymous):

see, Let S = y^2 =4ax be the equation of the parabola and P(x1,y1) be any pt. in the region of the parabola then S1= y^2 = 4ax if S1>0 then p is outside the parabola if S1 = 0 then P lies on the parabola if s1<0 then p lies inside the parabola

OpenStudy (anonymous):

Ok Ill try that.

OpenStudy (anonymous):

you'll surely get the answer :)